Stony Brook Update T.K. Hemmick for the Tent Crew
Getting back in business… • We have completed the dry run of the QE system. • Doing the dry run was a good idea since several small snags were encountered and fixed: • Evaporator system had errant sensor that claimed cooling water was not flowing. • pAmmeter had broken wire inside…ruined PMT reference measurement until repair. • Now under vacuum and functioning perfectly. 4.3 x 10-8 mbar
Result from Monday • The result from Monday for the QE scan is shown in the figure. • The GEM used wasa veteran: • Extra GEM from final shot for east arm. • Used in gain test box for months. • Transported from GTBstorage and back. • Friday test shows excellent QE (like brand new), HOWEVER!!!
Friday looked different. • Friday, the PMT circuitry was not working resulting in the PMT having no bias. • However, current from the cathode was measured, nonetheless. • The Friday cathode current was only ~70% of the Monday cathode current. • Scenarios: • Lamp was low and we had no reference…OK. • QE was low and recovered with a weekend of heavy pumping…not OK.
Plan for today: • Last night we prepared 3 GEMs from the HBD for their QE test…chosen principally by flash lamp test result. (good&bad). • We chose to leave these GEMs in the transport box in the glovebox until today. • This will allow us to get a first QE measure only an hour after initial pumpdown. • If things look other than normal, we’ll wait under high vacuum and test again.
Other News • LEGS gas system is at Stony Brook. • It is big. • It needs modifications: • Increase flow capacity to meet needs. • Decrease certain system impedances. • Build protection circuitry: • Protect HBD during underpressure. • Protect HBD during overpressure. • Protect HBD during power outage (& restore). • Jack Mahon is working with us to make all the necessary changes.